The following T-accounts represent September activity.

Required:

Compute the missing amounts indicated by the letters (a) through (i).

Materials Inventory
BB (9/1) 8,500
(a) 3,700
(b)
EB (9/30) 9,300
Work-In-Process Inventory
BB (9/1) 24,200
178,600 (e)
121,000
90,500
EB (9/30) 16,100
Finished Goods Inventory
BB (9/1) 14,500
(e) (f)
EB (9/30) (g)
Cost of Goods Sold
396,000
Applied Overhead Control
(d)
Manufacturing Overhead Control
121,000
3,700
36,200
32,600
3,700
Wages Payable
124,300
162,000 (c)
36,200
119,500 EB (9/30)
Accumulated Depreciation—Plant & Equipment
202,400 BB (9/1)
(h)
235,000 EB (9/30)
Accounts Payable—Material Suppliers
95,000
Prepaid Expenses
BB(9/1) 24,200
(i)
EB(9/30) 20,500
Answer #1

(b) $178600

(a) 178600+3700+9300-8500 = $183100

(c) $121000

(d) $90500

(e) 24200+178600+121000+90500-16100 = $398200

(f) $396000

(g) $14500+398200-396000 = $16700

(h) $235000 - 202400 = $32600

(i) 24200-20500 = $3700

If you have any query, kindly comment with your query and please mark thumbs up.